### v1

@@ IV-A Numerical Results for APCC-SIMPLE: Attacks Without Knowledge
In Table I and Table II, we show the numerical results for various cases we have tested. The numerical results illustrate that for these cases where we apply the APCC-SIMPLE approach with optimized MLE fusion to address attackers without protection system knowledge, the NRMSE between the APCC-SIMPLE (SIMPLE in Tables) estimated trajectory and the true trajectory (last column) is small and it is very close (identical to two decimal places) to that for the GE (using MLE). It indicates the APCC-SIMPLE approach successfully identifies the attacked sensors and excludes their readings accordingly in the estimation/fusion process.
TABLE I: APCC-SIMPLE Gaussian Results $\beta=99.9\%$
$\sigma^{2}$ is noise variance, $N$ is ##\## of sensors, $N_{A}$ is ##\## of attacked sensors, NRMSE GE and NRMSE SIMPLE (APCC-SIMPLE) both use MLE
$\sigma^{2}$ $N$ $N_{A}$ NRMSE GE NRMSE SIMPLE
@@
TABLE II: APCC-SIMPLE Laplacian Results $\beta=99.9\%$
See TABLE I for column definitions.
$\sigma^{2}=2b^{2}$ $N$ $N_{A}$ NRMSE GE NRMSE SIMPLE
@@
IV-B Numerical Results for APCC-ADDITIONAL and Attacks With Knowledge
IV-B1 Edge Attack - APCC-SIMPLE vs APCC-ADDITIONAL
@@
TABLE III: Gaussian Non-Random Edge Attack $\alpha=90\%$ $\beta=99.9\%$
@@
TABLE IV: Gaussian Random Edge Attack $\alpha=90\%$ $\beta=99.9\%$
@@
For the case when the sensor noise is Gaussian, Table III (Non-Random attack) and Table IV (Random attack) show the NRMSE performance improvements of the APCC-ADDITIONAL approach over the APCC-SIMPLE approach for the edge attacks. For large $N_{A}$, the APCC-ADDITIONAL approach nearly perfectly identifies and removes all the edge attacked sensors to provide NRMSE performance close to the GE (identical to two decimal places). The performance of the APCC-ADDITIONAL approach is considerably better than that of the performance of the APCC-SIMPLE approach for the edge attacks which is extremely important in some critical applications. However, considering its simplicity. the performance of the APCC-SIMPLE approach for the edge attacks is not that bad and might be suitable in some noncritical applications. Note that the degradation of the worst-case attack (Edge attack) is bounded to a reasonably small value for APCC-SIMPLE but this will not be the case if we remove the APCC-SIMPLE processing as discussed previously. In the cases considered, the nonrandom attacks are more powerful than the corresponding random attacks as expected. Similar conclusions to those drawn from Table III and Table IV can be drawn from the cases shown in Table V (NonRandom attack) and Table VI (Random attack) which consider cases with Laplacian noise as opposed to Gaussian noise.
TABLE V: Laplacian Non-Random Edge Attack $\alpha=90\%$ $\beta=99.9\%$
@@
TABLE VI: Laplacian Random Edge Attack $\alpha=90\%$ $\beta=99.9\%$
@@
TABLE VII: Gaussian Optimized Non-Random Water Filling Attack
@@
NRMSE for $\alpha=80\%$, $\beta=80\%$ $4.15\times 10^{-5}$ $1.29\times 10^{-4}$ $2.56\times 10^{-4}$ $4.03\times 10^{-4}$ $4.47\times 10^{-4}$
@@
NRMSE for $\alpha=50\%$, $\beta=80\%$ $6.06\times 10^{-5}$ $1.22\times 10^{-4}$ $1.40\times 10^{-4}$ $1.40\times 10^{-4}$ $1.40\times 10^{-4}$
@@
IV-B2 Water-filling Attack - Worst-case attack on APCC-ADDITIONAL processing
The previously considered edge attack is the worst-case (causes the most damage) attack for the APCC-SIMPLE processing, but the worst-case attack for the APCC-ADDITIONAL processing is the water-filling attack we discussed previously. In the water-filling attack considered in this subsection, we optimally select the sensors to attack, so we call this a nonrandom attack. Table VII illustrates the APCC-ADDITIONAL processing performance (NRMSE with MLE fusion) for several values of $\alpha$ and $\beta$ for the nonrandom water-filling attack for some cases with Gaussian sensor noise. Note that while the majority of the results in Table VII are for water-filling attacks (worst-case for ADDITIONAL), the first row in Table VII below the headings is for edge attacks (worst-case for SIMPLE) to allow simple comparison.
The results in Table VII show that, with proper choice of the parameters $\alpha$ and $\beta$, we can find approaches that make the worst-case NRMSE performance of the APCC-ADDITIONAL processing with a large $N_{A}$ smaller than that of the worst-case NRMSE performance of APCC-SIMPLE processing in the same case and close to GE. As Table VII illustrates, this decrease in worst-case NRMSE performance of the APCC-ADDITIONAL processing for cases with a large $N_{A}$ comes with a small increase in the NRMSE for cases with $N_{A}=0$ when compared with the APCC-SIMPLE approach in Table III (same as GE). Table VII shows that when the noise distribution is Gaussian with a variance of $1\times 10^{-4}$, using $\alpha=80\%$ and $\beta=80\%$ yields worst-case performance with $N_{A}=40$ that is within a factor of $2$ of the GE for that case while yielding performance with $N_{A}=0$ that is within a factor of $1.8$ of the GE for that case. While this seems to provide good performance at either extreme, other choices of $\alpha$ and $\beta$ allow different tradeoffs.
Table VIII illustrates similar findings for cases with Laplacian noise. Figure 8 illustrates the trade-offs that the different previously considered $(\alpha,\beta)$ approaches can achieve in terms of worst-case NRMSE performance ($y$-axis) for a large $N_{A}$ versus the $N_{A}=0$ NRMSE performance ($x$-axis) under Gaussian sensor noise. In addition, Figure 9 shows the trade-offs in worst-case NRMSE performance ($y$-axis) for a large $N_{A}$ compared to the no-attack NRMSE performance ($x$-axis) when subjected to Laplacian sensor noise.
It’s important to note that any error in the prediction in Fig. 3 could negatively affect the performance of the proposed APCC approach. We omit results showing this since the idea is quite intuitive. An accurate prediction approach must be employed. One might be concerned that there could be an accumulation of prediction errors over time. Thus if some attacked sensors are determined to be unattacked at a given time step, these sensors will be used in the prediction for the next time step as described in Fig 3, possibly increasing the prediction error at the next step. This error could accumulate over time if the attacks always move the prediction in a consistent direction. However, experimental results show if we employ a good machine learning approach which is properly trained then by properly adjusting $\beta$ and $\alpha$ in APCC-ADDITIONAL, the impact of even powerful optimized attacks always moving the prediction in a consistent direction do not cause large degradation over longer trajectories. Table IX shows results when the noise distribution is Gaussian with a variance of $1\times 10^{-4}$ for an attacker launching 40 optimum (non-random water-filling, consistent direction over time) attacks each time step on a total of 50 sensors. Table IX show that by setting $\alpha$ to 60% and $\beta$ to 80%, the APCC-ADDITIONAL approach provides enough protection to counteract even this severe worst-case attack. The simulation results show that the NRMSE reaches $1.86\times 10^{-4}$ after 150 time steps. Even when extending the simulation to $7.0\times 10^{4}$ time steps, the NRMSE only increases to $2.80\times 10^{-4}$, about 1.5 times larger. We observed similarly small increases for slightly larger $\alpha$ of $70\%$ or $80\%$. These results demonstrate that with proper protection, the damage from attacks leaking through and degrading the prediction can be made small, even under the worst-case attack, allowing for acceptable performance.
TABLE VIII: Laplacian Optimized Non-Random Water Filling Attack
@@ IV-B2 Water-filling Attack - Worst-case attack on APCC-ADDITIONAL processing
NRMSE for $\alpha=40\%$, $\beta=60\%$ $8.99\times 10^{-5}$ $1.18\times 10^{-4}$ $1.18\times 10^{-4}$ $1.18\times 10^{-4}$
@@
NRMSE for $\alpha=60\%$, $\beta=70\%$ $4.70\times 10^{-4}$ $1.99\times 10^{-3}$ $1.99\times 10^{-3}$ $1.99\times 10^{-3}$
@@
NRMSE for $\alpha=40\%$, $\beta=70\%$ $6.60\times 10^{-4}$ $1.13\times 10^{-3}$ $1.13\times 10^{-3}$ $1.13\times 10^{-3}$
@@
TABLE IX: Comparison between NRMSEs for short run and long run
V Conclusions And Future Work

### v2

@@ 4.1 Numerical Results for APCC-SIMPLE: Attacks Without Knowledge
In Table 1 and Table 2, we show the numerical results for various cases we have tested. The numerical results illustrate that for these cases where we apply the APCC-SIMPLE approach with optimized MLE fusion to address attackers without protection system knowledge, the NRMSE between the APCC-SIMPLE (SIMPLE in Tables) estimated trajectory and the true trajectory (last column) is small and it is very close (identical to two decimal places) to that for the GE (using MLE). It indicates the APCC-SIMPLE approach successfully identifies the attacked sensors and excludes their readings accordingly in the estimation/fusion process.
Table 1: APCC-SIMPLE Gaussian Results $\beta=99.9\%$
$\sigma^{2}$ is noise variance, $N$ is #\# of sensors, $N_{A}$ is #\# of attacked sensors, NRMSE GE and NRMSE SIMPLE (APCC-SIMPLE) both use MLE. $m=150.$
$\sigma^{2}$ NN $N_{A}$ NRMSE GE NRMSE SIMPLE
@@
Table 2: APCC-SIMPLE Laplacian Results $\beta=99.9\%$
See TABLE 1 for column definitions. $m=150.$
$\sigma^{2}=2b^{2}$ NN $N_{A}$ NRMSE GE NRMSE SIMPLE
@@
Table 3: = APCC-SIMPLE Gaussian with $m=4.8\times 10^{5}$.
See TABLE I for column definitions.
$\sigma^{2}$ NN $N_{A}$ NRMSE GE NRMSE SIMPLE
$1.0\times 10^{-4}$ 50 40 $5.19\times 10^{-5}$ $5.19\times 10^{-5}$
$1.0\times 10^{-4}$ 50 30 $3.67\times 10^{-5}$ $3.67\times 10^{-5}$
$1.0\times 10^{-4}$ 50 10 $2.60\times 10^{-5}$ $2.60\times 10^{-5}$
$1.0\times 10^{-4}$ 20 10 $5.19\times 10^{-5}$ $5.19\times 10^{-5}$
$1.0\times 10^{-4}$ 10 5 $7.34\times 10^{-5}$ $7.34\times 10^{-5}$
$1.0\times 10^{-2}$ 50 40 $5.19\times 10^{-4}$ $5.19\times 10^{-4}$
$1.0\times 10^{-2}$ 50 30 $3.67\times 10^{-4}$ $3.67\times 10^{-4}$
$1.0\times 10^{-2}$ 50 10 $2.60\times 10^{-4}$ $2.60\times 10^{-4}$
$1.0\times 10^{-2}$ 20 10 $5.19\times 10^{-4}$ $5.19\times 10^{-4}$
$1.0\times 10^{-2}$ 10 5 $7.34\times 10^{-4}$ $7.34\times 10^{-4}$
Table 4: APCC-SIMPLE Laplacian with $m=4.8\times 10^{5}$.
See TABLE I for column definitions.
$\sigma^{2}=2b^{2}$ NN $N_{A}$ NRMSE GE NRMSE SIMPLE
$2.0\times 10^{-4}$ 50 40 $6.25\times 10^{-5}$ $6.25\times 10^{-5}$
$2.0\times 10^{-4}$ 50 30 $4.24\times 10^{-5}$ $4.24\times 10^{-5}$
$2.0\times 10^{-4}$ 50 10 $2.89\times 10^{-5}$ $2.89\times 10^{-5}$
$2.0\times 10^{-4}$ 20 10 $6.25\times 10^{-5}$ $6.25\times 10^{-5}$
$2.0\times 10^{-4}$ 10 5 $9.73\times 10^{-5}$ $9.73\times 10^{-5}$
$2.0\times 10^{-2}$ 50 40 $6.25\times 10^{-4}$ $6.25\times 10^{-4}$
$2.0\times 10^{-2}$ 50 30 $4.24\times 10^{-4}$ $4.24\times 10^{-4}$
$2.0\times 10^{-2}$ 50 10 $2.89\times 10^{-4}$ $2.89\times 10^{-4}$
$2.0\times 10^{-2}$ 20 10 $6.25\times 10^{-4}$ $6.25\times 10^{-4}$
$2.0\times 10^{-2}$ 10 5 $9.73\times 10^{-4}$ $9.73\times 10^{-4}$
In Table 3, we performed exactly the same experiment as in Table 1, but used a true trajectory (similar to Figure 7(a)) that is longer, with $m=4.8\times 10^{5}$ instead of $m=150$. Since the true trajectory is different, as expected from (3), the normalization factor in the NRMSE changed. This change makes the GE slightly larger than in Table 1, but the change is just as expected given the new true trajectory. According to the results shown in Table 3, our APCC-SIMPLE approach yields NRMSE values identical to those of GE with two decimal places in all tested cases. For example, when 40 out of 50 sensors are under attack and the variance of the Gaussian noise is $1.0\times 10^{-4}$, the NRMSE of the APCC-SIMPLE approach is $5.19\times 10^{-5}$, which matches the NRMSE of the GE in this new case to two decimal places, which proves our APCC-SIMPLE approach works well in this case.
Similar results for Laplacian noise are shown in Table 4, which performs exactly the same experiment as conducted in Table 3 but changes the noise distribution from Gaussian to Laplacian. According to the results shown in Table 4, our APCC-SIMPLE approach produces NRMSE values identical to those of GE to two decimal places in all tested cases. For instance, when 40 out of 50 sensors are under attack and the variance of the Laplacian noise is $2.0\times 10^{-4}$, the NRMSE of the APCC-SIMPLE approach is $6.25\times 10^{-5}$, which matches the NRMSE of the GE in this new case to two decimal places. Thus our APCC-SIMPLE approach works well in this case also.
In the next few results to follow, we go back to using $m=150$ unless otherwise stated. We defined $m$ as the trajectory length. Why don’t you use this? Making $m$ longer is easy to explain. Refer back to (2). In Table 3 we conduct the same experiment as in Table 1, but using a true waveform of much longer duration while keeping otherwise identical settings. The numerical results in Table 3 confirm that our APCC-SIMPLE approach performs reliably under this extended scenario. In particular, the NRMSE of GE remains consistently close to that of APCC-SIMPLE, demonstrating that the proposed method maintains robust performance even for long waveforms.
4.2 Numerical Results for APCC-ADDITIONAL and Attacks With Knowledge
4.2.1 Edge Attack - APCC-SIMPLE vs APCC-ADDITIONAL
@@
Table 5: Gaussian Non-Random Edge Attack $\alpha=90\%$ $\beta=99.9\%$ $m=150.$
@@
Table 6: Gaussian Random Edge Attack $\alpha=90\%$ $\beta=99.9\%$ $m=150.$
@@
For the case when the sensor noise is Gaussian, Table 5 (Non-Random attack) and Table 6 (Random attack) show the NRMSE performance improvements of the APCC-ADDITIONAL approach over the APCC-SIMPLE approach for the edge attacks. For large $N_{A}$, the APCC-ADDITIONAL approach nearly perfectly identifies and removes all the edge attacked sensors to provide NRMSE performance close to the GE (identical to two decimal places). The performance of the APCC-ADDITIONAL approach is considerably better than that of the performance of the APCC-SIMPLE approach for the edge attacks which is extremely important in some critical applications. However, considering its simplicity. the performance of the APCC-SIMPLE approach for the edge attacks is not that bad and might be suitable in some noncritical applications. Note that the degradation of the worst-case attack (Edge attack) is bounded to a reasonably small value for APCC-SIMPLE but this will not be the case if we remove the APCC-SIMPLE processing as discussed previously. In the cases considered, the nonrandom attacks are more powerful than the corresponding random attacks as expected. Similar conclusions to those drawn from Table 5 and Table 6 can be drawn from the cases shown in Table 7 (NonRandom attack) and Table 8 (Random attack) which consider cases with Laplacian noise as opposed to Gaussian noise.
Table 7: Laplacian Non-Random Edge Attack $\alpha=90\%$ $\beta=99.9\%$ $m=150.$
@@
Table 8: Laplacian Random Edge Attack $\alpha=90\%$ $\beta=99.9\%$ $m=150.$
@@
Table 9: Gaussian Optimized Non-Random Water Filling Attack. $m=150.$
@@
NRMSE for $\alpha=80\%$, $\beta=80\%$ $4.15\times 10^{-5}$ $1.29\times 10^{-4}$ $2.56\times 10^{-4}$ $03\times 10^{-4} 4.4$$47\times 10^{-4}$
@@
NRMSE for $\alpha=50\%$, $\beta=80\%$ $6.06\times 10^{-5}$ $1.22\times 10^{-4}$ $\times 10^{-4} 1.40×104$$\times 10^{-4}$ $1.40\times 10^{-4}$
@@
Table 10: Laplacian Optimized Non-Random Water Filling Attack. $m=150.$
@@
NRMSE for $\alpha=40\%$, $\beta=60\%$ $8.99\times 10^{-5}$ $18\times 10^{-4} 1.1$$8\times 10^{-4} 1.1$$18\times 10^{-4}$
@@
NRMSE for $\alpha=60\%$, $\beta=70\%$ $4.70\times 10^{-4}$ $9\times 10^{-3} 1.99×1$$9\times 10^{-3}$ $1.99\times 10^{-3}$
@@
NRMSE for $\alpha=40\%$, $\beta=70\%$ $6.60\times 10^{-4}$ $13\times 10^{-3} 1.1$$3\times 10^{-3} 1.1$$13\times 10^{-3}$
@@
Table 11: Comparison between NRMSEs for additional processing with $m=150,m=70,000,$ and $m=300,000$.
@@
NRMSE (300,000 Steps) $6.00\times 10^{-4}$
4.2.2 Water-filling Attack - Worst-case attack on APCC-ADDITIONAL processing
The previously considered edge attack is the worst-case (causes the most damage) attack for the APCC-SIMPLE processing, but the worst-case attack for the APCC-ADDITIONAL processing is the water-filling attack we discussed previously. In the water-filling attack considered in this subsection, we optimally select the sensors to attack, so we call this a nonrandom attack. Table 9 illustrates the APCC-ADDITIONAL processing performance (NRMSE with MLE fusion) for several values of $\alpha$ and $\beta$ for the nonrandom water-filling attack for some cases with Gaussian sensor noise. Note that while the majority of the results in Table 9 are for water-filling attacks (worst-case for ADDITIONAL), the first row in Table 9 below the headings is for edge attacks (worst-case for SIMPLE) to allow simple comparison.
The results in Table 9 show that, with proper choice of the parameters $\alpha$ and $\beta$, we can find approaches that make the worst-case NRMSE performance of the APCC-ADDITIONAL processing with a large $N_{A}$ smaller than that of the worst-case NRMSE performance of APCC-SIMPLE processing in the same case and close to GE. As Table 9 illustrates, this decrease in worst-case NRMSE performance of the APCC-ADDITIONAL processing for cases with a large $N_{A}$ comes with a small increase in the NRMSE for cases with $N_{A}=0$ when compared with the APCC-SIMPLE approach in Table 5 (same as GE). Table 9 shows that when the noise distribution is Gaussian with a variance of $1\times 10^{-4}$, using $\alpha=80\%$ and $\beta=80\%$ yields worst-case performance with $N_{A}=40$ that is within a factor of $2$ of the GE for that case while yielding performance with $N_{A}=0$ that is within a factor of $1.8$ of the GE for that case. While this seems to provide good performance at either extreme, other choices of $\alpha$ and $\beta$ allow different tradeoffs.
Table 10 illustrates similar findings for cases with Laplacian noise. Figure 8 illustrates the trade-offs that the different previously considered $(\alpha,\beta)$ approaches can achieve in terms of worst-case NRMSE performance ($y$-axis) for a large $N_{A}$ versus the $N_{A}=0$ NRMSE performance ($x$-axis) under Gaussian sensor noise. In addition, Figure 9 shows the trade-offs in worst-case NRMSE performance ($y$-axis) for a large $N_{A}$ compared to the no-attack NRMSE performance ($x$-axis) when subjected to Laplacian sensor noise.
It’s important to note that any error in the prediction in Fig. 3 could negatively affect the performance of the proposed APCC approach. We omit results showing this since the idea is quite intuitive. An accurate prediction approach must be employed. One might be concerned that there could be an accumulation of prediction errors over time. Thus if some attacked sensors are determined to be unattacked at a given time step, these sensors will be used in the prediction for the next time step as described in Fig 3, possibly increasing the prediction error at the next step. This error could accumulate over time if the attacks always move the prediction in a consistent direction. However, experimental results show if we employ a good machine learning approach which is properly trained then by properly adjusting $\beta$ and $\alpha$ in APCC-ADDITIONAL, the impact of even powerful optimized attacks always moving the prediction in a consistent direction do not cause large degradation over longer trajectories. Table 11 shows results when the noise distribution is Gaussian with a variance of $1\times 10^{-4}$ for an attacker launching 40 optimum (non-random water-filling, consistent direction over time) attacks each time step on a total of 50 sensors. Table 11 show that by setting $\alpha$ to 60% and $\beta$ to 80%, the APCC-ADDITIONAL approach provides enough protection to counteract even this severe worst-case attack. The simulation results show that the NRMSE reaches $1.86\times 10^{-4}$ after $m=150$ time steps. Even when increasing $m$ to $m=7.0\times 10^{4}$ and $m=3.0\times 10^{5}$ time steps, the NRMSE increases only slightly, to $2.80\times 10^{-4}$ and $6.00\times 10^{-4}$, corresponding to approximately $1.5\times$ and $3.2\times$ the original value, respectively. These results demonstrate that with proper protection, the damage from attacks leaking through and degrading the prediction can be made small, even under the worst-case attack, allowing for acceptable performance.
4.3 Comparisons with Subspace-based Sparse Attack Detection (SSAD) Approach
Table 12: SSAD for Gaussian noise with specific shift of $1.0\times 10^{4}$. See TABLE I for column definitions. $m=150$.
$\sigma^{2}$ NN $N_{A}$ NRMSE SIMPLE NRMSE SSAD
$1.0\times 10^{-4}$ 50 40 $5.03\times 10^{-5}$ $1.22\times 10^{+2}$
$1.0\times 10^{-4}$ 50 30 $3.56\times 10^{-5}$ $9.23\times 10^{+1}$
$1.0\times 10^{-4}$ 50 10 $2.52\times 10^{-5}$ $3.10\times 10^{+1}$
$1.0\times 10^{-4}$ 20 10 $5.03\times 10^{-5}$ $7.87\times 10^{+1}$
$1.0\times 10^{-4}$ 10 5 $12\times 10^{-5} 7.7$$77\times 10^{+1}$
$1.0\times 10^{-2}$ 50 40 $5.03\times 10^{-4}$ $1.22\times 10^{+2}$
$1.0\times 10^{-2}$ 50 30 $3.56\times 10^{-4}$ $9.23\times 10^{+1}$
$1.0\times 10^{-2}$ 50 10 $2.52\times 10^{-4}$ $3.10\times 10^{+1}$
$1.0\times 10^{-2}$ 20 10 $5.03\times 10^{-4}$ $7.87\times 10^{+1}$
$1.0\times 10^{-2}$ 10 5 $12\times 10^{-4} 7.7$$77\times 10^{+1}$
In Table 12 for $m=150$, we compare the performance of the subspace-based sparse attack detection approach (SSAD) described in [42] with that of our APCC-SIMPLE approach. In Table 12 we always launch an attack that shifts the actual trajectory at attacked sensors by adding $1.0\times 10^{4}$. The details of the number of sensors attacked, total number of sensors and the variance of the Gaussian noise added are given in ($N,N_{A},\sigma^{2}$) in Table 12. The particular sensors attacked are chosen randomly. Consider the case where 40 out of 50 sensors are under attack and the variance of the Gaussian noise is $1.0\times 10^{-4}$. The NRMSE of the APCC-SIMPLE approach is $5.03\times 10^{-5}$, which is close to the NRMSE of GE, while the NRMSE of SSAD is $1.22\times 10^{+2}$, which is much larger. According to the results in Table 12, our APCC-SIMPLE approach consistently yields much lower NRMSEs in all the cases tested in the numerical experiment.
To be fair, the attack detection approach in [42] was shown to perform well for sparse attacks in [42], for which it is designed, but not for the non-sparse attacks. This explains Table 12. Note that there are many other approaches, as discussed in the literature review, which are designed for cases where less than $k$ sensors are attacked (or similar assumptions), However, these methods tend to perform very poorly for other cases (more than $k$ sensors attacked) and all cases are allowed by our attack model.
4.4 Comparison with Data-Reconstruction (DR) Approach
Table 13: DR for Gaussian noise with specific shift of $1.0\times 10^{4}$. See TABLE I for column definitions. $m=150$.
$\sigma^{2}$ NN $N_{A}$ NRMSE GE NRMSE RECON
$1.0\times 10^{-4}$ 50 40 $5.03\times 10^{-5}$ $1.26\times 10^{2}$
$1.0\times 10^{-4}$ 50 30 $3.56\times 10^{-5}$ $9.48\times 10^{1}$
$1.0\times 10^{-2}$ 50 40 $5.03\times 10^{-4}$ $1.26\times 10^{2}$
$1.0\times 10^{-2}$ 50 30 $3.56\times 10^{-4}$ $9.48\times 10^{1}$
In Table 13 for $m=150$, we compare the performance of the data-reconstruction (DR) approach from [43] with that of our APCC-SIMPLE method. The total number of sensors $N$, number of attacked sensors $N_{A}$, and the Gaussian noise variance $\sigma^{2}$ are listed in Table 13, and the attack procedure is described in Subsection 4.3.
Consider the representative case with $N=50$, $N_{A}=40$, and $\sigma^{2}=1.0\times 10^{-4}$. The NRMSE of APCC-SIMPLE is $5.03\times 10^{-5}$, which is close to the GE, whereas the NRMSE of the DR method is $1.26\times 10^{2}$, several orders of magnitude larger. Across all tested cases, APCC-SIMPLE yields substantially lower NRMSE and is always close to the GE.
To be fair, the attack detection approach in [43] was shown to perform well for certain assumptions on the number of attacked sensors. These assumptions are not true for all attacks in our attack model, which explains the bad behavior.
4.5 Comparisons with EDAD Approaches
As previously discussed, one can always substitute any valid trajectory for another valid trajectory and the substituted trajectory will pass EDAD since it is valid. Such substitutions are possible, as long as the shape of the trajectory after attack is similar to what is expected based on the training data. There are always cases where these substitutions can lead to very large errors as we try to explain here. Let’s focus on a sensor measuring position since we focus on this in the numerical examples. Let’s also focus on the case where the actual trajectory $y_{1},y_{2},\ldots,y_{m}$ in the computation of (3) satisfies ${\frac{1}{m}\sum_{j=1}^{m}|y_{j}|}=1$ so we can show bad performance for EDAD for some cases in a simple way. Thus, an easy way to change the position trajectory to another valid trajectory $\hat{y}_{1},\hat{y}_{2},\ldots,\hat{y}_{m}$, with similar shape, is to shift all the elements of the position trajectory by some constant $h$, thus $\hat{y}_{j}=y_{j}+h$ for all $1\leq j\leq m$. The large values of $h$ that are possible in typical applications lead to large NRMSEs in (3). For example, a shift so that $\hat{y}_{j}=y_{j}+h$ for all $1\leq j\leq m$, yields $NRMSE=h$. Thus if $h$ is unbounded, so is $NRMSE$. Even if you limit the trajectories to come from a specified area, the errors can be very large. One can estimate the diameter of Bethlehem PA as approximately 3970 meters. Thus if we set $h=3970$, then $NRMSE=3970$ which is much larger than even the large errors in Table 12. The point is that attacks which pass EDAD can lead to extremely large errors so one can’t employ this approach since you can’t typically tolerate these large errors.
Note that both of our approaches build on the simple approach. The simple approach limits the size by which the attack can change the sensor data. If the sensor data is changed by too large an amount, due to the confidence interval comparison in Fig 3, the attack will be blocked. On the other hand, large attacks can pass EDAD if they follow valid trajectories. To pass our approaches, the attack can’t be too different from the ”typical unattacked sensor prediction” in Fig 3. How large an attack is allowed is controlled by $\beta$. The size allowed is very small in all the cases we considered, on the order of $\beta\%$ of the sensor noise in the training data. Thus our approaches limit the size of attacks, while EDAD does not.
5 Conclusions And Future Work